# Enhancing Logical Expressiveness in Graph Neural Networks via Path-Neighbor Aggregation

### v2

@@ Introduction
A Knowledge Graph (KG) organizes and represents knowledge in a structured form. In a KG, the nodes represent entities (such as people, places, or concepts), while the edges capture relationships between entities (such as “belongs to” or “contains”). KG reasoning typically requires models to understand and infer complex relations within graph structures. In recent years, graph neural networks (GNNs) (wu2020comprehensive; velivckovic2017graph; hamilton2017inductive; xu2018powerful) have shown great potential in KG reasoning tasks due to their powerful ability to represent graph-structured data. GNNs can automatically learn intricate relations between nodes from data and effectively capture both global and local structural information via message passing mechanisms, thereby substantially enhancing reasoning capabilities.
@@
From the perspective of the logical expression ability of GNN, relational aggregation methods such as R-GCN (schlichtkrull2018modeling) and CompGCN (vashishth2019composition), cannot perform rule learning well. The GNN method that dynamically adjusts based on conditions, such as NBFNet (zhu2021neural), RED-GNN (zhang2022knowledge), and A*Net (zhu2024net), has been proven to have strong logical expressiveness (understanding; ATheory) and can learn basic logical relationships, such as unary prediction, negation, conjunction, disjunction, etc. Based on the Relational Weisfeiler-Leman (R-WL) (Barcelo00O22) test, the limitations of the logical rule learning ability of existing relational GNN (R-GNN) models were clarified, and it was proved that these models cannot capture higher-order logical rules (ATheory). In other words, it was theoretically proved that the rule expression ability of conditional GNN (C-GNN) is strictly more powerful than that of R-GNN method. Although the labeling trick can enhance the rule learning ability of C-GNN in KG link prediction by introducing specific labels to nodes, it still faces two limitations: on the one hand, it will reduce the generalization ability of the model; on the other hand, the design of adding constant labels makes it difficult to apply to KG reasoning tasks in inductive settings.
@@
For example, in Figure 2, for the rule $C_{3}(u,v):=\exists\,z_{1},z_{2}(r_{1}(u,z_{1})\wedge r_{2}(z_{1},z_{2})\wedge r_{3}(z_{2},v))$, when $z_{1}$ and $z_{2}$ in the rule are variables, it can match any element that satisfies the variable constraint; but once the variable is fixed to a constant c, that is, it becomes $C_{3}^{\prime}(u,v):=\exists\ z_{1}(r_{1}(u,z_{1})\wedge r_{2}(z_{1},\textit{c})\wedge r_{3}(\textit{c},v))$, only the specific elements corresponding to c can satisfy the rule, so the scope of application of the rule is limited, and the coverage (Quinlan1990; MUGGLETON1994629) naturally becomes smaller, thereby weakening the generalization ability of the model.
@@
The expressive power of GNNs is often analyzed through graph isomorphism testing. Vanilla GNNs have been shown to match the 1-WL test (xu2018powerful), with extensions to KGs (Barcelo00O22). To enhance expressivity, two main directions have emerged: designing stronger WL variants (morris2019weisfeiler; otto2023graded; Barcelo00O22), and applying initialization techniques such as node labeling or identity features (AbboudCGL21; you2021identity; SatoYK21). In link prediction, the structural expressivity of GNN alone has been found to be insufficient (Srinivasan020), and the incorporation of labeling enables GNNs to learn structural information, although the existing results are primarily limited to single relational graphs (LabelingTrick). Beyond WL-based approaches, GNNs have been shown to model first-order logic (FOL) with two variables ($\text{FOC}_{2}$), and their expressivity in node classification can be described using the counting extension of modal logic (barcelo2020logical; cai1992optimal). This has been extended to KGs, linking logical and WL-based expressivity and identifying the rule structures that GNNs can learn (ATheory). Inspired by Teru, Denis, and Hamilton (2020), further improvements combine conditional message passing with labeling to enhance the learning of logical rules in KG reasoning.
@@
GNN-based KG reasoning can be categorized into R-GNN and C-GNN (ATheory). R-GNN methods like R-GCN and CompGCN perform relation-aware message passing to update entity embeddings, which are then used to score candidate links between entities.
C-GNN methods generate query-aware representations by marking the head entity and propagating messages accordingly, and can be categorized into subgraph-based and progressive aggregation methods. Subgraph-based models, such as GraIL (teru2020inductive) and CoMPILE (mai2021communicative), extract local neighborhoods between entity pairs and apply GNNs to propagate messages within subgraphs. Progressive aggregation methods, such as NBFNet (zhu2021neural), RED-GNN (zhang2022knowledge), and A*Net (zhu2024net), further enhance reasoning capabilities. NBFNet incorporates the Bellman-Ford (goldberg1993heuristic) to propagate messages along multi-hop paths. RED-GNN determines the next hop based on the current node, the query relation, and the next relation embedding. A*Net introduces A* search to enhance inference efficiently.
@@
Graph neural networks based on conditional message passing mechanisms (zhu2021neural; zhang2022knowledge) have a common message passing mechanism. Given a triple query $(u,q,?$), where u is the head entity and q is the query relation. Conditional Graph Neural Networks (C-GNN) first mark the head entity u with the query relation q to distinguish it from all other nodes in the KG. Then, it iteratively calculates and updates the representations of entity pairs ${\textbf{h}}_{v|u,q}$ of all entities in the KG that can reach the tail v from the head u:
@@
The logical expressivity of a GNN refers to the set of logical formulas it can learn (understanding). C-GNN demonstrates superior logical expressive power compared to R-GNN, leading to enhanced performance (ATheory; understanding). We use the counting extension of modal logic (CML), which extends the counting expressiveness of graded modal logic, to describe FOL properties in KG. CML consists of propositional constants (always true $\top$, always false $\bot$), unary logical predicates $P_{i}(x)$, and recursive rules. Given FOL $\varphi(x,y)$, $\psi(x,y)$, positive integer variables $N\geq 1$ and $r\in R$, according to the recursive rules, $\neg\varphi(x,y)$, $\varphi(x,y)\wedge\psi(x,y)$, and $\exists^{N\geq n}z(\varphi(x,z)\wedge r(z,y))$ are also in CML. Based on the CML in KG, the logical rule expressive power of C-GNN can be expressed as follows:
@@
Chain-like rules $C_{3}$, as illustrated in Figure 3(a), represent a commonly used rule structure in KG reasoning methods that rely on paths or rules (understanding; ATheory). The corresponding rule formula is expressed as follows:
@@
For example, in Figure 3, U and T cannot be effectively distinguished. When a constant $c$ is added to $z_{1}$ and a specific initial representation is given, the C-GNN can learn the structure of the rules $U^{\prime}(u,v):=r_{1}(u,c)\wedge(\exists\,z_{2},z_{3}(r_{2}(c,z_{2})\wedge r_{3}(z_{2},v)\wedge r_{4}(c,z_{3})\wedge r_{5}(z_{3},v))$. EL-GNN (understanding) improves the rule learning ability by using the labeling trick based on C-GNN. However, when nodes are given unique labels, the model may over-rely on this “artificial” information during training, thereby ignoring other potential topological features or relationship clues, resulting in a decrease in the model’s generalization ability.
Next, we introduce the concept of “coverage” (Quinlan1990; MUGGLETON1994629) to assess the logical expressive power of these rules and explain why assigning constants to variables reduces coverage.
@@
Here, $(u_{i}^{{}^{\prime}},q,v_{i}^{{}^{\prime}})$ denotes a negative sample. Negative samples are constructed under the Partial Completeness Assumption (PCA) (bordes2013translating; sun2019rotate) by randomly replacing the head (or tail) entity with an incorrect one from the knowledge graph, and $n$ denotes the number of negative samples. In order to balance logical expressive power and efficiency, we use a 2-hop PN-GNN.
@@
Results on synthetic datasets are shown in Table  1. The results based on the R-GNN method come from Qiu et al. (2024). The results of the EL-GNN method are obtained by using NBFNet as the backbone and carefully searching for parameters on the validation set using the Ray (liaw2018tune). Through the analysis of the results, we can observe that:
@@
The results of the transductive setting are shown in Table  2. Across all evaluation metrics, C-GNN based methods surpass those based on representation learning, paths, and R-GNNs. This outcome aligns with C-GNN theory, showing that C-GNN can learn and leverage logical rule structures. They also indicate that, although GNNs capture more graph structures than path-based methods, the ability to learn logical rule structures is crucial for KG reasoning. PN-GNN and EL-GNN further improve upon C-GNN, demonstrating that enhancing the representation of logical rule structures leads to better inference performance. On the real-world dataset FB15K237, which contains $U$ and $T$ structures, EL-GNN and PN-GNN outperform C-GNN. Notably, the node out-degree $d$ is a sensitive hyperparameter for EL-GNN and has a great impact on performance (understanding), whereas PN-GNN does not require this parameter.
@@ Acknowledgements
This work is particularly supported by the National Key Research and Development Program of China (No. 2022YFB3104103), the National Natural Science Foundation of China (No. 62302507), and the Provincial Natural Science Foundation of Hunan (No. 2024JJ7249, 2025JJ70302).
@@ A.1 The logical Expressiveness of C-GNN
As a supplement to the main text, this section first introduces the formal definition of CML in KG (ATheory; understanding) to facilitate the proofs, and then presents the formal proofs of Theorems 1 and 2, which characterize the logical expressiveness of C-GNN.
@@
Theorems 1 and 2 are adapted from Theorem 3.2 in (understanding) and Theorem 5.3 in (ATheory). For implementation-level details, we refer the reader to the original sources. We restate these theorems here for convenience:
@@
Assume for a contradiction that there exists a C-GNN that can learn $\varphi(x)$. Since $\varphi(x)$ is not equivalent to any formula in CML, according to Theorem C.7 in (understanding) and Theorem 2.2 in (otto2023graded), there exists two KGs $G$ and $G^{\prime}$ and two entities $v$ in $G$ and $v^{\prime}$ in $G^{\prime}$ such that $\mathrm{Unr}^{L}_{G}(v)\cong\mathrm{Unr}^{L}_{G^{\prime}}(v^{\prime})$ for every $L\in\mathbb{N}$ and such that $G,v\models\varphi$ and $G^{\prime},v^{\prime}\not\models\varphi$, where $\mathrm{Unr}^{L}_{G}(v)$ represents unraveling tree, for a detailed definition refer to (understanding; ATheory). Furthermore, according to the definition of RWL-test in (Barcelo00O22), because $\mathrm{Unr}^{L}_{G}(v)\cong\mathrm{Unr}^{L}_{G^{\prime}}(v^{\prime})$ for every $L\in\mathbb{N}$, we have $e_{v}^{(L)}=e_{v^{\prime}}^{(L)}$. So, if a formula $\varphi(x)$ is not equivalent to any formula in CML, there is no C-GNN can learn $\varphi(x)$.
@@ B.1 Datasets
The synthetic datasets $C_{3}$, $C_{4}$, $I_{1}$, $I_{2}$, $T$, and $U$ are adopted from (understanding), while we additionally construct $T_{label}$ and $U_{label}$ to study the effect of labeling strategies. The process of generating $T_{label}$ and $U_{label}$ consists of three steps, namely, generation of the logical rule structure, generation of noisy triplets, and completion of missing triplets.
  • (1) Following the steps described in (understanding), we construct the training set and the validation set.
@@
Transductive setting means that the test set does not contain the target triples in the training set, but all entities and relations are allowed to appear in the training set. We selected two widely used datasets WN18RR (dettmers2018convolutional) and FB15K237 (toutanova2015representing) to verify the effectiveness of the model. The statistical information of the datasets is shown in Table B.2.
@@
Inductive setting refers to the entities in the training set and the test set that do not overlap, but all relations in the test set have appeared in the training set. The widely used WN18RR (v1-v4) and FB15K237 (v1-v4) datasets in the inductive setting were constructed by Komal et al. (teru2020inductive). The statistics are shown in Table B.3.
@@ B.2 Hyperparameter Settings
In the experimental setup in synthetic datasets, the hyperparameter search space is defined as follows: the learning rate is selected from $[0.0001,0.001,0.01,0.1]$, embedding dimensions from $[32,64,128,256,512]$, number of GNN layers from $[4,5,6,7]$, batch size from $[20,40,60]$ and dropout rate within the range $(0,0.6)$. The activation function is chosen from (ReLU, Tanh, IDD, Sigmoid), the aggregation function from $\{\text{{sum, min, max}}\}$, and the score function adopts either TranE (bordes2013translating) or DistMult (yang2014embedding).
For C-GNN and PN-GNN, we employ Ray (liaw2018tune) to perform automatic hyperparameter tuning on the valid. PN-GNN integrates path-neighbors via concatenation to enhance node representations, while C-GNN utilizes sum aggregation.
In the transductive setting, the experimental configurations for FB15K237 are as follows: the embedding dimension is set to 64, the number of GNN layers is 6, the relational message passing function adopts DistMult, and the aggregation function uses PNA (corso2020principal). The number of negative samples is 128, meaning each positive sample corresponds to 128 negative samples. The optimizer is Adam (kingma2014adam), with a learning rate of 0.005 and a batch size of 32. Additionally, we apply LayerNorm (ba2016layer) during training to stabilize the process. For the WN18RR, the embedding dimension remains 64, the number of GNN layers is 6, and the relational message passing function employs RotatE (sun2019rotate). The batch size is set to 16, and the aggregation function remains PNA. The number of negative samples is 32, with the optimizer and learning rate settings kept consistent with FB15K237.

### v3

@@ Introduction
A Knowledge Graph (KG) organizes and represents knowledge in a structured form. In a KG, the nodes represent entities (such as people, places, or concepts), while the edges capture relationships between entities (such as “belongs to” or “contains”). KG reasoning typically requires models to understand and infer complex relations within graph structures. In recent years, graph neural networks (GNNs) (Wu et al. 2020; Velickovic et al. 2018; Hamilton et al. 2017; Xu et al. 2019) have shown great potential in KG reasoning tasks due to their powerful ability to represent graph-structured data. GNNs can automatically learn intricate relations between nodes from data and effectively capture both global and local structural information via message passing mechanisms, thereby substantially enhancing reasoning capabilities.
@@
From the perspective of the logical expression ability of GNN, relational aggregation methods such as R-GCN (Schlichtkrull et al. 2018) and CompGCN (Vashishth et al. 2020), cannot perform rule learning well. The GNN method that dynamically adjusts based on conditions, such as NBFNet (Zhu et al. 2021), RED-GNN (Zhang and Yao 2022), and A*Net (Zhu et al. 2023), has been proven to have strong logical expressiveness (Qiu et al. 2024; Huang et al. 2023) and can learn basic logical relationships, such as unary prediction, negation, conjunction, disjunction, etc. Based on the Relational Weisfeiler-Leman (R-WL) (Barceló et al. 2022) test, the limitations of the logical rule learning ability of existing relational GNN (R-GNN) models were clarified, and it was proved that these models cannot capture higher-order logical rules (Huang et al. 2023). In other words, it was theoretically proved that the rule expression ability of conditional GNN (C-GNN) is strictly more powerful than that of R-GNN method. Although the labeling trick can enhance the rule learning ability of C-GNN in KG link prediction by introducing specific labels to nodes, it still faces two limitations: on the one hand, it will reduce the generalization ability of the model; on the other hand, the design of adding constant labels makes it difficult to apply to KG reasoning tasks in inductive settings.
@@
For example, in Figure 2, for the rule $C_{3}(u,v):=\exists\,z_{1},z_{2}(r_{1}(u,z_{1})\wedge r_{2}(z_{1},z_{2})\wedge r_{3}(z_{2},v))$, when $z_{1}$ and $z_{2}$ in the rule are variables, it can match any element that satisfies the variable constraint; but once the variable is fixed to a constant c, that is, it becomes $C_{3}^{\prime}(u,v):=\exists\ z_{1}(r_{1}(u,z_{1})\wedge r_{2}(z_{1},\textit{c})\wedge r_{3}(\textit{c},v))$, only the specific elements corresponding to c can satisfy the rule, so the scope of application of the rule is limited, and the coverage (Quinlan 1990; Muggleton and de Raedt 1994) naturally becomes smaller, thereby weakening the generalization ability of the model.
@@
The expressive power of GNNs is often analyzed through graph isomorphism testing. Vanilla GNNs have been shown to match the 1-WL test (Xu et al. 2019), with extensions to KGs (Barceló et al. 2022). To enhance expressivity, two main directions have emerged: designing stronger WL variants (Morris et al. 2019; Otto 2023; Barceló et al. 2022), and applying initialization techniques such as node labeling or identity features (Abboud et al. 2021; You et al. 2021; Sato et al. 2021). In link prediction, the structural expressivity of GNN alone has been found to be insufficient (Srinivasan and Ribeiro 2020), and the incorporation of labeling enables GNNs to learn structural information, although the existing results are primarily limited to single relational graphs (Zhang et al. 2021). Beyond WL-based approaches, GNNs have been shown to model first-order logic (FOL) with two variables ($\text{FOC}_{2}$), and their expressivity in node classification can be described using the counting extension of modal logic (Barceló et al. 2020; Cai et al. 1992). This has been extended to KGs, linking logical and WL-based expressivity and identifying the rule structures that GNNs can learn (Huang et al. 2023). Inspired by Teru, Denis, and Hamilton (2020), further improvements combine conditional message passing with labeling to enhance the learning of logical rules in KG reasoning.
@@
GNN-based KG reasoning can be categorized into R-GNN and C-GNN (Huang et al. 2023). R-GNN methods like R-GCN and CompGCN perform relation-aware message passing to update entity embeddings, which are then used to score candidate links between entities.
C-GNN methods generate query-aware representations by marking the head entity and propagating messages accordingly, and can be categorized into subgraph-based and progressive aggregation methods. Subgraph-based models, such as GraIL (Teru et al. 2020) and CoMPILE (Mai et al. 2021), extract local neighborhoods between entity pairs and apply GNNs to propagate messages within subgraphs. Progressive aggregation methods, such as NBFNet (Zhu et al. 2021), RED-GNN (Zhang and Yao 2022), and A*Net (Zhu et al. 2023), further enhance reasoning capabilities. NBFNet incorporates the Bellman-Ford (Goldberg and Radzik 1993) to propagate messages along multi-hop paths. RED-GNN determines the next hop based on the current node, the query relation, and the next relation embedding. A*Net introduces A* search to enhance inference efficiently.
@@
Graph neural networks based on conditional message passing mechanisms (Zhu et al. 2021; Zhang and Yao 2022) have a common message passing mechanism. Given a triple query $(u,q,?$), where u is the head entity and q is the query relation. Conditional Graph Neural Networks (C-GNN) first mark the head entity u with the query relation q to distinguish it from all other nodes in the KG. Then, it iteratively calculates and updates the representations of entity pairs ${\textbf{h}}_{v|u,q}$ of all entities in the KG that can reach the tail v from the head u:
@@
The logical expressivity of a GNN refers to the set of logical formulas it can learn (Qiu et al. 2024). C-GNN demonstrates superior logical expressive power compared to R-GNN, leading to enhanced performance (Huang et al. 2023; Qiu et al. 2024). We use the counting extension of modal logic (CML), which extends the counting expressiveness of graded modal logic, to describe FOL properties in KG. CML consists of propositional constants (always true $\top$, always false $\bot$), unary logical predicates $P_{i}(x)$, and recursive rules. Given FOL $\varphi(x,y)$, $\psi(x,y)$, positive integer variables $N\geq 1$ and $r\in R$, according to the recursive rules, $\neg\varphi(x,y)$, $\varphi(x,y)\wedge\psi(x,y)$, and $\exists^{N\geq n}z(\varphi(x,z)\wedge r(z,y))$ are also in CML. Based on the CML in KG, the logical rule expressive power of C-GNN can be expressed as follows:
@@
Chain-like rules $C_{3}$, as illustrated in Figure 3(a), represent a commonly used rule structure in KG reasoning methods that rely on paths or rules (Qiu et al. 2024; Huang et al. 2023). The corresponding rule formula is expressed as follows:
@@
For example, in Figure 3, U and T cannot be effectively distinguished. When a constant $c$ is added to $z_{1}$ and a specific initial representation is given, the C-GNN can learn the structure of the rules $U^{\prime}(u,v):=r_{1}(u,c)\wedge(\exists\,z_{2},z_{3}(r_{2}(c,z_{2})\wedge r_{3}(z_{2},v)\wedge r_{4}(c,z_{3})\wedge r_{5}(z_{3},v))$. EL-GNN (Qiu et al. 2024) improves the rule learning ability by using the labeling trick based on C-GNN. However, when nodes are given unique labels, the model may over-rely on this “artificial” information during training, thereby ignoring other potential topological features or relationship clues, resulting in a decrease in the model’s generalization ability.
Next, we introduce the concept of “coverage” (Quinlan 1990; Muggleton and de Raedt 1994) to assess the logical expressive power of these rules and explain why assigning constants to variables reduces coverage.
@@
Here, $(u_{i}^{{}^{\prime}},q,v_{i}^{{}^{\prime}})$ denotes a negative sample. Negative samples are constructed under the Partial Completeness Assumption (PCA) (Bordes et al. 2013; Sun et al. 2019) by randomly replacing the head (or tail) entity with an incorrect one from the knowledge graph, and $n$ denotes the number of negative samples. In order to balance logical expressive power and efficiency, we use a 2-hop PN-GNN.
@@
Results on synthetic datasets are shown in Table  1. The results based on the R-GNN method come from Qiu et al. (2024). The results of the EL-GNN method are obtained by using NBFNet as the backbone and carefully searching for parameters on the validation set using the Ray (Liaw et al. 2018). Through the analysis of the results, we can observe that:
@@
The results of the transductive setting are shown in Table  2. Across all evaluation metrics, C-GNN based methods surpass those based on representation learning, paths, and R-GNNs. This outcome aligns with C-GNN theory, showing that C-GNN can learn and leverage logical rule structures. They also indicate that, although GNNs capture more graph structures than path-based methods, the ability to learn logical rule structures is crucial for KG reasoning. PN-GNN and EL-GNN further improve upon C-GNN, demonstrating that enhancing the representation of logical rule structures leads to better inference performance. On the real-world dataset FB15K237, which contains $U$ and $T$ structures, EL-GNN and PN-GNN outperform C-GNN. Notably, the node out-degree $d$ is a sensitive hyperparameter for EL-GNN and has a great impact on performance (Qiu et al. 2024), whereas PN-GNN does not require this parameter.
@@ Acknowledgements
This work is particularly supported by the National Key Research and Development Program of China (No. 2022YFB3104103), the National Natural Science Foundation of China (No. 62302507), the Provincial Natural Science Foundation of Hunan (No. 2024JJ7249, 2025JJ70302), and the National First-Class Undergraduate Major (Network Engineering) Fund.
@@ A.1 The logical Expressiveness of C-GNN
As a supplement to the main text, this section first introduces the formal definition of CML in KG (Huang et al. 2023; Qiu et al. 2024) to facilitate the proofs, and then presents the formal proofs of Theorems 1 and 2, which characterize the logical expressiveness of C-GNN.
@@
Theorems 1 and 2 are adapted from Theorem 3.2 in (Qiu et al. 2024) and Theorem 5.3 in (Huang et al. 2023). For implementation-level details, we refer the reader to the original sources. We restate these theorems here for convenience:
@@
Assume for a contradiction that there exists a C-GNN that can learn $\varphi(x)$. Since $\varphi(x)$ is not equivalent to any formula in CML, according to Theorem C.7 in (Qiu et al. 2024) and Theorem 2.2 in (Otto 2023), there exists two KGs $G$ and $G^{\prime}$ and two entities $v$ in $G$ and $v^{\prime}$ in $G^{\prime}$ such that $\mathrm{Unr}^{L}_{G}(v)\cong\mathrm{Unr}^{L}_{G^{\prime}}(v^{\prime})$ for every $L\in\mathbb{N}$ and such that $G,v\models\varphi$ and $G^{\prime},v^{\prime}\not\models\varphi$, where $\mathrm{Unr}^{L}_{G}(v)$ represents unraveling tree, for a detailed definition refer to (Qiu et al. 2024; Huang et al. 2023). Furthermore, according to the definition of RWL-test in (Barceló et al. 2022), because $\mathrm{Unr}^{L}_{G}(v)\cong\mathrm{Unr}^{L}_{G^{\prime}}(v^{\prime})$ for every $L\in\mathbb{N}$, we have $e_{v}^{(L)}=e_{v^{\prime}}^{(L)}$. So, if a formula $\varphi(x)$ is not equivalent to any formula in CML, there is no C-GNN can learn $\varphi(x)$.
@@ B.1 Datasets
The synthetic datasets $C_{3}$, $C_{4}$, $I_{1}$, $I_{2}$, $T$, and $U$ are adopted from (Qiu et al. 2024), while we additionally construct $T_{label}$ and $U_{label}$ to study the effect of labeling strategies. The process of generating $T_{label}$ and $U_{label}$ consists of three steps, namely, generation of the logical rule structure, generation of noisy triplets, and completion of missing triplets.
  • (1) Following the steps described in (Qiu et al. 2024), we construct the training set and the validation set.
@@
Transductive setting means that the test set does not contain the target triples in the training set, but all entities and relations are allowed to appear in the training set. We selected two widely used datasets WN18RR (Dettmers et al. 2018) and FB15K237 (Toutanova et al. 2015) to verify the effectiveness of the model. The statistical information of the datasets is shown in Table B.2.
@@
Inductive setting refers to the entities in the training set and the test set that do not overlap, but all relations in the test set have appeared in the training set. The widely used WN18RR (v1-v4) and FB15K237 (v1-v4) datasets in the inductive setting were constructed by Komal et al. (Teru et al. 2020). The statistics are shown in Table B.3.
@@ B.2 Hyperparameter Settings
In the experimental setup in synthetic datasets, the hyperparameter search space is defined as follows: the learning rate is selected from $[0.0001,0.001,0.01,0.1]$, embedding dimensions from $[32,64,128,256,512]$, number of GNN layers from $[4,5,6,7]$, batch size from $[20,40,60]$ and dropout rate within the range $(0,0.6)$. The activation function is chosen from (ReLU, Tanh, IDD, Sigmoid), the aggregation function from $\{\text{{sum, min, max}}\}$, and the score function adopts either TranE (Bordes et al. 2013) or DistMult (Yang et al. 2015).
For C-GNN and PN-GNN, we employ Ray (Liaw et al. 2018) to perform automatic hyperparameter tuning on the valid. PN-GNN integrates path-neighbors via concatenation to enhance node representations, while C-GNN utilizes sum aggregation.
In the transductive setting, the experimental configurations for FB15K237 are as follows: the embedding dimension is set to 64, the number of GNN layers is 6, the relational message passing function adopts DistMult, and the aggregation function uses PNA (Corso et al. 2020). The number of negative samples is 128, meaning each positive sample corresponds to 128 negative samples. The optimizer is Adam (Kingma and Ba 2015), with a learning rate of 0.005 and a batch size of 32. Additionally, we apply LayerNorm (Ba et al. 2016) during training to stabilize the process. For the WN18RR, the embedding dimension remains 64, the number of GNN layers is 6, and the relational message passing function employs RotatE (Sun et al. 2019). The batch size is set to 16, and the aggregation function remains PNA. The number of negative samples is 32, with the optimizer and learning rate settings kept consistent with FB15K237.